Created by Main.MatteoSolfaroli on 2007-02-12 - 14:44
Main activities ongoing (check meeting minutes for further details):
- LHCCOOL cards needed to access sector 78
- Cooldown of Sec78 from 80K to 20K planned for this week
- Cryo-leak being repaired in magnet 3002. The problem appeared also in other magnets (SSS349).
- Powering Procedure Test from Wednesday 14th
- Warm Magnets powering tests ongoing this week in LSS8L an LSS8R

RAT_20070212_WithPPT_planning.pdf:
https://twiki.cern.ch/twiki/pub/HCC/BlogEntryPoint8x2007x02x12x14x24/RAT_20070212_WithPPT_planning.pdf
Edit, attach file or add comment